OpenStudy (anonymous):

1)which equation represents a circle whose center is (3 -2) 2)which graph represents a circle with the equation (x-5)2+(y+1)2=9

OpenStudy (hunus):

The equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\] Where h and k are the coordinates (h, k) of the center of the circle and r is the radius of the circle. With the (3, -2) you're given what is the equation of the circle?

OpenStudy (anonymous):

there is four choices are (x+3)2+(y-2)2=4

OpenStudy (hunus):

(3, -2) -> (h, k) So h = 3 k = -2 Try putting these values in the equation I gave you